[Date Prev][Date Next][Thread Prev][Thread Next][Date Index][Thread Index]
Everything seemed to fall apart last night. I did not watch the desiccant
indicator and TOM1 iced up. The RA drive on TOM2 was stripped. No telling
what I did wrong. Then TOM2 and TOM3 cameras got out of focus. We put in
a new cable for the TOM3 output and it appears to be faulty. We also
worked on the shutter for TOM2 I camera and this appears to be fixed. You
win some of em.
So it was a busy evening. I got everything back up and running by 10PM or
so. This except for TOM1 with was so iced up that it did not recover until
near dawn. Then TOM1 had some temperature control problem that seems to
have now gone away. Just a typical evening on the production line.
By 7:30 AM or so I had all the data there was recorded on 9 CDs. A total
of about 600 images. I then started processing. Everything that could
make it through the pipeline was done by noon. No TOM1 data made it due to
the temperature control problem on the V camera. We got about 260,000 V, I
pairs from TOM2 and TOM3. For newcomers to the list, this means a few
thousand measurements will turn out to be from variable stars. Many will
be yet to be identified variables.
Meanwhile I am near the end of processing the old data. There are about 20
CD's with a total of 1000 images left to process. Then I will be current
and hope to get results within 24 hours of processing. I can imagine
getting the data to the data base next day now that I have an ADSL
line. Not so fast, but a lot faster than a modem. A good days data will
be 50 Mbytes, so it will take a while.
It looks like I will have about 16 million total measurement pairs. Most
of them between -6 and +18. This should give us most of the variables
between mag 8 and 13 in this region.